# ComputerRL: Scaling End-to-End Online Reinforcement Learning for Computer Use Agents

### v1

@@ 3.2 Reinforcement Learning from Verifiable Rewards
Reward Design. We employ a rule-based reward framework to furnish transparent and verifiable training signals in the context of reinforcement learning. For each trajectory generated by the agent, task completion is assessed via an automated verification function. If the trajectory successfully solves the task, we assign a reward of 1 to every action that is both correctly formatted and substantially contributes to the solution; otherwise, any trajectory failing the verification or actions with improper formatting receive a reward of 0. Distinct from conventional approaches that propagate step-wise returns via the Bellman equation (Sutton et al., 1998) and infer advantages retrospectively, our methodology treats each prompt-response pair as an independent training instance, with rewards allocated according to the final trajectory outcome. This direct, step-level reward assignment provides explicit and interpretable feedback by tightly coupling agent behaviors with task success, thereby facilitating more efficient and robust policy optimization during reinforcement learning.
@@ 3.3 Entropulse for Scaling RL Training
In the RL training phase described in Section 3.2, we observe that model performance plateaus after hundreds of training steps, as indicated by stagnating task completion rates and decreasing entropy. This premature convergence motivates us to investigate strategies for extending effective training and enhancing policy exploration. Inspired by DAPO (Yu et al., 2025), we experiment with increasing the clipping threshold. While this adjustment attenuates the decline in entropy, it also significantly slows the rate of policy improvement.
To address these limitations, we propose Entropulse, an approach motivated by the observation that the objectives of SFT and RL differ markedly during training. Specifically, as entropy decreases during RL optimization, integrating SFT at critical junctures enhances the agent’s exploration and trajectory diversity, thus facilitating further policy optimization. In practice, during the initial RL stage, we aggregate and retain all successful rollout trajectories generated during training runs. Conventionally, these trajectories are utilized only once and subsequently discarded, resulting in substantial underutilization of computational resources. Our collected dataset contains trajectories from various policies at different training stages, representing a valuable and often overlooked source of diverse behavioral data.
We process this dataset by randomly selecting successful trajectories per unique task to construct a new SFT training set, which exhibits the following salient attributes:
@@
    Diversity: Rollouts originate from heterogeneous policies and training states, offering a variety of problem-solving strategies.
@@
SFT training on this curated dataset produces notable shifts in policy behavior. While evaluation task performance remains stable, the resulting policy shows increased entropy relative to the original one, indicating enhanced exploration. Building upon this enhanced exploration capability, we conducted a second round of RL training, which yielded significant performance improvements and enabled us to achieve state-of-the-art results on the OSWorld benchmark.
@@ 4 Experiments
We employ ComputerRL on two open-source base LLMs, GLM-4-9B-0414 and Qwen2.5-14B, to produce AutoGLM-OS-9b/14b, as part of the AutoGLM project Liu et al. (2024b). For experiment, we conduct extensive experiments across various scenarios to evaluate the performance of the AutoGLM-OS in executing a range of tasks within the Ubuntu desktop environment compared to common baselines. To closely reflect the real user experience, we select OSWorld (Xie et al., 2024) that utilizes interactive environments for evaluation.
Table 1: AutoGLM-OS performance on OSWorld and OSWorld-Verified (updated in 2025.08). We compare AutoGLM-OS with state-of-the-art agent models, including both proprietary and open models.
@@ 4.1 Main Results
We train AutoGLM-OS with GLM-4-9B-0414 and Qwen2.5-14B as the backbone and evaluate on the OSWorld (Xie et al., 2024) and OSWorld-Verified benchmark  [^3] https://xlang.ai/blog/osworld-verified, comparing its performance against state-of-the-art models, including CUA (OpenAI, 2025), Claude-4 (Anthropic, 2023), UI-TARS (Qin et al., 2025), and Agent S2 (Agashe et al., 2025), among others. The comparative results are in Table 1.
The results indicate that AutoGLM-OS achieves superior performance across a range of domains, with its advantages most pronounced in the challenging multi-apps setting. Notably, AutoGLM-OS exhibits exceptional proficiency in long-range planning and complex reasoning tasks. Moreover, by employing the API-GUI strategy, AutoGLM-OS can accomplish tasks using at most 1/3 of the steps required by the strongest baseline approaches, thereby demonstrating remarkable gains in execution efficiency. These results collectively underscore the potential of ComputerRL to advance the state of the art in automated task completion across heterogeneous application environments.
@@ 4.2 Ablation Study
To evaluate the influence of various algorithms and training datasets on agent performance, we present an ablation study in Table 2.
@@
Framework Ablation. We compare the performance of the GUI-only approach with our proposed API-GUI paradigm, using GPT-4o as the base model. The results demonstrate that the API-GUI paradigm substantially outperforms the GUI-only baseline across all domains. Specifically, the API-GUI strategy achieves an average success rate of 26.2%, representing a 134% improvement over the GUI-only approach (11.2%). The most significant gains are observed in the Office (27.9% vs. 6.2%) and Professional (41.6% vs. 14.3%) domains, where API-GUI provides 350% and 191% improvements, respectively. These results validate our core hypothesis that combining API calls with GUI interactions enables more efficient and reliable task execution, particularly for complex professional workflows that benefit from programmatic control.
Training Ablation. In the training ablation, we study the progressive impact of different training stages using Qwen2.5-14B as the base model. Starting from the baseline, we observe consistent performance improvements through each training phase. Behavior Cloning (BC) establishes a solid foundation with an average success rate of 31.9%. The first RL phase (RL1) yields substantial gains, increasing the average performance to 42.0% (+10.1%). Interestingly, Entropulse phase maintains similar performance at 41.5% while significantly increasing action entropy, which enhances exploration diversity and enables the final RL2 phase to achieve further improvements. The final RL2 phase achieves the best performance at 45.8% (+3.8% from RL1), benefiting from the increased exploration capacity introduced by Entropulse. Notably, the Workflow domain shows the most dramatic improvement throughout training (10.8% → 27.2%), while the other domains maintain consistently high performance. These results underscore the importance of multi-stage training, particularly the effectiveness of RL in fine-tuning agent behaviors for executing complex tasks.
Figure 5: ComputerRL training curves of reward (left) and entropy (right) with 95% confidence intervals. The red line denotes the training with entropy recovery via Entropulse after the initial RL stage, while the grey line denotes continued training with only reference resetting.
@@
The results demonstrate that incorporating Entropulse increases the model’s entropy, thereby restoring its exploratory capacity. This enhanced exploration substantially scales the effective training steps and ultimately leads to improved overall performance.
@@ 4.3 Case Study and Error Analysis
We conduct a case study in the Ubuntu environment to identify potential avenues for system optimization. Although our model exhibits robust performance across most scenarios, several limitations have been identified. In particular, errors encountered during task execution can be categorized into four primary types: visual perception errors, multi-application coordination failures, operational illusions, and other errors. The distribution of these error types is presented in Table 3.
@@
Additional representative examples, including both successful and unsuccessful cases, are provided in Appendix A to illustrate the model’s capabilities and limitations further.
Table 3: Error distribution of AutoGLM-OS.
Error Type Proportion
Vision 25.8%
Multiple App 34.4%
Illusion 14.2%
Others 25.6%
5 Future Direction

### v2

@@ 3.2 Reinforcement Learning with Verifiable Rewards
Reward Design. We select a subset of the constructed human-annotated data (in Section 3.1) for RL and employ a rule-based verification function to provide verifiable training signals for each trajectory. Successfully solved trajectories receive a reward of 1 for every correctly formatted action that contributes to the solution; failed trajectories or improperly formatted actions receive a reward of 0. Unlike conventional approaches that propagate step-wise returns via the Bellman equation, our methodology treats each prompt-response pair as an independent training instance with rewards based on the final trajectory outcome. This direct reward assignment provides explicit feedback by coupling agent behaviors with task success, facilitating effective policy optimization.
@@ 3.3 Entropulse for Scaling RL Training
In the RL training in Section 3.2, we observe that model performance plateaus after hundreds of training steps, with stagnating task completion rates and decreasing entropy. This premature convergence motivates us to investigate strategies for extending effective training and enhancing policy exploration. Inspired by DAPO (Yu et al., 2025), we experiment with increasing the clipping threshold, which attenuates the decline in entropy but significantly slows down policy improvement.
To address the issue, we propose Entropulse, motivated by the observation that SFT and RL objectives differ markedly during training. As entropy decreases during RL optimization, integrating SFT at critical junctures enhances exploration and trajectory diversity, facilitating further policy optimization. During initial RL training, we aggregate and retain all successful rollout trajectories. While conventionally discarded after single use, these trajectories from various policies at different training steps represent valuable and diverse behavioral data.
We process this dataset by randomly selecting successful trajectories per unique task to construct a new SFT training set, which exhibits the following attributes:
@@
    Diversity: Rollouts originate from heterogeneous policies in different training steps, offering a variety of problem-solving strategies.
@@
SFT training on this curated dataset produces notable shifts in policy behavior. While evaluation task performance remains stable, the resulting policy shows increased entropy relative to the original one, indicating enhanced exploration. Building upon this enhanced exploration capability, we conduct a second round of RL training, which yields significant performance improvements and enables us to achieve state-of-the-art results in computer automation. The training details are in Appendix D.
@@ 4 Experiments
We employ ComputerRL on GLM-4-9B-0414 (GLM et al., 2024) and GLM-4.1V-9B-Thinking (Hong et al., 2025), to produce AutoGLM-OS-9b. We conduct extensive experiments across various scenarios to evaluate AutoGLM-OS’s performance within the computer environment.
Table 1: AutoGLM-OS performance on OSWorld and OSWorld-Verified (updated in 2025.08). We compare AutoGLM-OS with state-of-the-art agents, including both proprietary and open models.
@@ 4.1 Main Results
To closely reflect the real user experience, we evaluate AutoGLM-OS on the OSWorld (Xie et al., 2024) and OSWorld-Verified benchmark, comparing its performance against state-of-the-art models, including CUA (OpenAI, 2025), Claude-4 (Anthropic, 2023), and UI-TARS (Qin et al., 2025), among others. The comparative results are in Table 1. The results indicate that AutoGLM-OS achieves superior performance across a range of domains, with its advantages most pronounced in the challenging multi-apps setting. Moreover, by employing the API-GUI strategy, AutoGLM-OS can accomplish tasks using at most 1/3 of the steps required by the strongest baseline approaches, demonstrating remarkable gains in execution efficiency. These results underscore the potential of ComputerRL to advance the state of the art in computer automation across various applications.
4.2 Office Application Performance
As a critical interface for delivering and presenting, office application constitutes an important testbed for evaluating computer use agents. To assess agent performance in this domain, we curate a set of 180 challenging tasks from three sources: SpreadsheetBench (Ma et al., 2024), PPTC (Guo et al., 2023), and in-house developed Writer domain tasks. These tasks are adapted as necessary to integrate them into the OSWorld framework. The resulting benchmark, termed OfficeWorld, enables systematic measurement of agent capabilities in office-oriented scenarios. The results are in Table 2.
Table 2: AutoGLM-OS performance on OfficeWorld compared to common baselines. We employ the same framework (with tools) and test settings to ensure a fair comparison.
Agent Model Word Excel PPT Average
DeepSeek-V3.1 (Liu et al., 2024a) 6.7 35.0 21.7 21.1
DeepSeek-R1 Guo et al. (2025) 13.3 36.7 18.3 22.8
Claude 3.7 Sonnet (Anthropic, 2023) 15.0 25.0 25.0 21.7
Claude 4.0 Sonnet (Anthropic, 2023) 18.3 35.0 20.0 24.4
Gemini-2.5-Pro (Team et al., 2023) 5.0 11.7 20.0 12.2
GPT-4o (Hurst et al., 2024) 18.3 21.7 8.3 16.1
GPT-4.1 (Achiam et al., 2023) 21.7 25.0 28.3 25.0
OpenAI o3 (Jaech et al., 2024) 23.3 36.7 41.7 33.9
ComputerRL (ours)
 w/ GLM-4-9B-0414 21.7 58.3 43.3 41.1
 w/ GLM-4.1V-9B-Thinking 30.0 58.3 41.7 43.3
@@ 4.3 Ablation Study
To evaluate the influence of various algorithms and training datasets on agent performance, we present an ablation study on the OSWorld benchmark in Table 3.
@@
Framework Ablation. We compare the performance of the GUI-only approach with our proposed API-GUI paradigm using GPT-4o. The results demonstrate that the API-GUI paradigm substantially outperforms the GUI-only baseline across all domains. Specifically, the API-GUI strategy achieves an average success rate of 26.2%, representing a 134% improvement over the GUI-only approach (11.2%). The most significant gains are observed in the Office (27.9% vs. 6.2%) and Professional (41.6% vs. 14.3%) domains, where API-GUI provides 350% and 191% improvements, respectively. These results validate our core hypothesis that combining API calls with GUI interactions enables more efficient and reliable task execution, particularly for complex professional workflows that benefit from programmatic control.
Figure 5: ComputerRL training curves of reward (left) and entropy (right) with 95% confidence intervals. The red line denotes the training with entropy recovery via Entropulse after the first RL stage, while the grey line denotes continued training with only reference resetting.
Training Ablation. We study the progressive impact of different training stages using Qwen2.5-14B. Starting from the backbone, Behavior Cloning (BC) establishes a solid foundation with 31.9%. The first RL phase (RL1) yields substantial gains, increasing the performance to 42.0% (+10.1%). Interestingly, Entropulse phase maintains similar performance (41.5%) while significantly increasing action entropy, which enhances exploration diversity and enables the final RL2 phase to achieve further improvements. The RL2 phase achieves the best performance at 45.8% (+3.8% from RL1), benefiting from the increased exploration capacity introduced by Entropulse. Notably, the Workflow domain shows the most dramatic improvement throughout training (10.8% → 27.2%), while the other domains maintain consistently high performance, highlighting the importance of multi-stage training.
@@
The results demonstrate that incorporating Entropulse increases the model’s entropy, thereby restoring its exploratory capacity. This enhanced exploration substantially scales the effective training steps, ultimately leading to improved overall performance.
@@ 4.4 Case Study and Error Analysis
We conduct a case study in the desktop environment to identify potential avenues for system optimization. Although our model exhibits robust performance across most scenarios, several limitations have been identified. In particular, errors encountered during task execution can be categorized into four primary types: visual perception errors, multi-application coordination failures, operational illusions, and other errors. The distribution of these error types is presented in Figure 6.
@@
Additional representative examples, including both successful and unsuccessful cases, are provided in Appendix G to further illustrate the model’s capabilities and limitations.
5 Related Work
Large Language Models. LLMs, such as GPT-4 (Achiam et al., 2023), Gemini (Team et al., 2023), Claude-3 (Anthropic, 2023), Llama (Touvron et al., 2023a), ChatGLM (Zeng et al., 2022; Du et al., 2022), Qwen (Bai et al., 2023a; Team, 2024), and Deepseek (Liu et al., 2024a; Guo et al., 2025), have demonstrated remarkable capabilities in knowledge representation and natural language understanding, leading to diverse downstream applications. Vision-Language Models (VLMs) (Hong et al., 2023; 2025; Bai et al., 2023b; Hurst et al., 2024) further extend LLMs to multimodal inputs, enabling joint reasoning over text and images for tasks such as visual question answering.
Computer Use Agents. CogAgent (Hong et al., 2023) introduces multimodal GUI understanding. AutoGLM (Liu et al., 2024c) decouples planning and grounding with online RL improvement. OS-Atlas (Wu et al., 2024b) proposes a foundational GUI action model. Aguvis (Xu et al., 2024) enables cross-platform interaction through visual training. PC-Agent-E (He et al., 2025) utilizes trajectory boosting for enhanced proficiency. UI-TARS (Qin et al., 2025) performs human-like GUI interactions from screenshots. Agent S2 (Agashe et al., 2025) integrates grounding with hierarchical reasoning. CUA (OpenAI, 2025) offers programmable desktop automation.
Computer Use Benchmarks. WebArena (Zhou et al., 2023) provides simulated websites for online interactions, but has limitations: discrepancies from real-world environments and a web-only focus. Similar issues exist in other web-focused benchmarks (Yao et al., 2022; Koh et al., 2024; Chezelles et al., 2024; Miyai et al., 2025). Software engineering benchmarks (Jimenez et al., 2023; Yang et al., 2024a; Li et al., 2024; Zan et al., 2025; Padigela et al., 2025) lack comprehensive desktop evaluation. OSWorld (Xie et al., 2024) addresses these gaps with 369 tasks with 134 evaluation functions. Windows Agent Arena (Bonatti et al., 2024) expands this with 150+ Windows-based tasks.
RL Training for LLMs. PPO (Schulman et al., 2017) addresses instability in policy gradients for RL training. GRPO (Guo et al., 2025) extends PPO with group sampling and removes value updates. DAPO (Yu et al., 2025) refines GRPO for improved stability. ARPO (Lu et al., 2025) adds replay buffers and task selection. ZeroGUI (Yang et al., 2025b) leverages VLMs for automated task generation and reward estimation. Verl (Sheng et al., 2025) presents hybrid programming for dynamic scheduling. AReaL (Fu et al., 2025) achieves asynchronous RL by decoupling model generation.
6 Conclusion
In this work, we present ComputerRL, a novel computer use agent framework that advances desktop automation through the integration of API-based and GUI actions, efficient data collection, and scalable RL training. Our experiments on OSWorld and OfficeWorld demonstrate superior reasoning, accuracy, and generalization compared to prior approaches. This work lays the groundwork for more capable and persistent autonomous computer use agents, highlighting promising directions for future research in intelligent human-computer interaction.
Appendix A API Development Workflow
In this section, we detail the methodology for leveraging LLMs to automate API construction. We propose a semi-automated workflow wherein users need only supply exemplar tasks performed within the target application; the LLM then autonomously generates both the necessary API code and corresponding test cases. The workflow comprises three primary stages: requirement analysis, API implementation, and test case generation.
Requirement Analysis During the requirements analysis phase, users provide a set of task examples related to the target application as input. The workflow leverages the LLM to analyze these task instances, extracting the essential functionalities required for task completion. It then compares these requirements against the existing API interface definitions to identify potential gaps. If uncovered functionalities are detected, the system automatically generates new API interfaces along with their corresponding parameter specifications.
API Implementation Upon obtaining the interface definitions, the workflow systematically iterates over each interface and its associated parameters. For each specification, it leverages the designated Python libraries of the target application to implement the corresponding API functionalities. Additionally, the workflow incorporates error-handling mechanisms and logging to facilitate human debugging and maintenance. This automated approach not only streamlines API development but also enhances consistency and reusability across different application contexts.
This automated methodology substantially lowers the manual effort required, enabling the creation of application-specific API sets with minimal human intervention. As a result, the barrier for users to develop APIs for diverse applications is significantly reduced. We have developed API sets for multiple commonly used default applications in Ubuntu and integrated them into our Ubuntu virtual machine environment.
Appendix B Action Space
Our action space for AutoGLM-OS is shown in Table 4, and our API number for each application is in Table 5.
Table 4: Action space of AutoGLM-OS
Function Description
open_app(app_name) Open specified application (e.g., Chrome, Terminal).
click(coordinates, num_clicks,button_type) Click at coordinates [x, y] with the specified mouse button and number of clicks.
type(coordinates, text, overwrite, enter) Type text at coordinates; optionally overwrite existing content and/or press Enter.
drag_and_drop(drag_from, drop_on) Drag from [x1, y1] and drop onto [x2, y2].
scroll(coordinates, direction) Scroll at coordinates in direction (up / down).
switch_window(window_id) Switch focus to the window with given ID.
hotkey(keys) Press a key combination (e.g., [ctrl, c]).
quote(content) Record content for memory.
wait() Pause execution temporarily.
exit(success) Terminate task with success (True) or failure (False).
Table 5: Statistics of the number of available APIs per application
Application Number of APIs
Code 12
Chrome 11
LibreOffice Calc 27
LibreOffice Impress 22
LibreOffice Writer 19
VLC 12
Total 103
Appendix C Prompt Formulation for AutoGLM-OS
The design of the observation space is pivotal, as it directly constrains the upper bound of the agent’s performance. In this section, we detail the integration of the API set with GUI operations, alongside the incorporation of contextual information from the desktop environment, to systematically construct both the agent’s observation space and action space. This unified framework ensures that the designed observation and action spaces capture the complexity of real-world tasks, providing a solid foundation for robust agent learning and generalization.
Action space formulation The integration of a large number of APIs with GUI operations, while ensuring effective agent interaction, remains a significant challenge. In practice, we mitigate this complexity by dynamically detecting the currently active application to infer potentially relevant APIs, thereby reducing the number of available APIs and lowering the agent’s adaptation overhead. Furthermore, we use Python classes and descriptive docstrings to delineate each operation type, ensuring they are clearly interpretable by most LLMs. This object-oriented strategy enhances the model’s understanding and precision in performing operations. These classes are provided to the agent via the system prompt, enabling interaction through Python function calls. This design facilitates rapid agent adaptation and efficient generalization of operations across diverse applications. Additionally, the agent’s output format is standardized in the system prompt, which encourages the agent to interleave reasoning and action execution. This approach promotes enhanced planning and reflective capabilities within the agent, thus improving its overall performance in complex task execution scenarios.
Observation formulation To facilitate the effective perception and manipulation of GUIs by the model, we leverage the Python Accessibility Toolkit Service Provider Interface (pyatspi) to extract comprehensive attributes of desktop elements systematically. Each GUI element encompasses the element’s semantic type, visible text content, precise screen coordinates, and spatial dimensions. This structured representation enables the LLM agent to parse, ground, and reason over the GUI in a manner analogous to human users.
We present the element format of the environment a11y tree in our observation space as follows:
tag text position (center x & y) size (w & h)
The tag is the XML tag of the element, such as div or button. The text is the text content of the element, which can be empty for elements that do not have text. The position is represented by the center coordinates (x, y) of the element, and the size is represented by its width (w) and height (h).
For the multimodal model, the a11y tree is removed from the input. Instead, we capture the GUI screenshot at a resolution of 1920 × 1080 pixels (1080p) and subsequently resize it to 1280 × 720 pixels (720p), which serves as the input representation of the desktop environment.
Beyond the extraction of individual GUI components, we augment the input space with rich contextual metadata to provide a holistic depiction of the agent’s operational environment. Specifically, we provide a comprehensive enumeration of open desktop windows, including their hierarchical relationships, as well as the name and additional information of the currently focused application. To promote consistent and adaptive behavior, we also deliver feedback from the most recent GUI action or tool call, which may include environmental status updates, confirmations, or error signals.
The app format of the observation space is as follows:
Window ID App Name Title
The Window ID is the unique identifier of the application window, App Name is the name of the application, and Title is the title of the application window.
History formulation Given the extensive length of GUI observations and the inherent constraints imposed by the model’s context window, it is necessary to efficiently manage the input history across multiple interaction rounds. For each interaction, we omit redundant and detailed interface information while preserving the complete sequence of the agent’s reasoning process, actions taken, and the corresponding operation feedback. This approach ensures the retention of the essential operational trajectory, thereby maximizing the informativeness of the historical context while maintaining compatibility with the model’s capacity limitations.
Collectively, the components above constitute the observation space and action space of our agent. This representation not only enhances the agent’s environmental cognition but also enables better strategies for long-horizon planning and reasoning. As a result, the agent is better equipped to execute complex, multi-step tasks across diverse applications in the desktop environment.
Below is our detailed prompt organization for AutoGLM-OS:
You are an agent which follow my instruction and perform desktop computer tasks as instructed.
You have good knowledge of computer and good internet connection and assume your code will run on a computer for controlling the mouse and keyboard.
For each step, you will get an observation of the desktop by 1) screenshot; 2) current application name; 3) accessibility tree, which is based on AT-SPI library; 4) application info; 5) last action result.
You should first generate a plan for completing the task, confirm the previous results, reflect on the current status, then generate operations to complete the task in python-style pseudo code using the predefined functions.
Your output should STRICTLY follow the format:
<think>
{**YOUR-PLAN-AND-THINKING**}
</think>
‘‘‘python
{**ONE-LINE-OF-CODE**}
You will be provided access to the following methods to interact with the UI:
1. class Agent, a grounding agent which provides basic action space to interact with desktop.
2. class {tool_class_name}, which provides tools to interact with the current application {app_name}.
Here are the defination of the classes:
‘‘‘python
{class_content}
* Note:
- Your code should be wrapped in ‘‘‘python‘‘‘, and your plan and thinking should be wrapped in <think></think>.
- Only **ONE-LINE-OF-CODE** at a time.
- Each code block is context independent, and variables from the previous round cannot be used in the next round.
- Do not put anything other than python code in ‘‘‘python‘‘‘.
- You **can only use the above methods to interact with the UI**, do not invent new methods.
- Return with Agent.exit(success=True)‘ immediately after the task is completed.
- If you think cannot complete the task, **DO NOT keep repeating actions, just return with Agent.exit(success=False)‘.**
- The computers environment is Linux, e.g., Desktop path is ’/home/user/Desktop
- My computers password is password’, feel free to use it when you need sudo rights
[…]
breaklines=true,
[…]
$\begin{lstlisting}$
<|user|>
**Environment State (Omitted)**
<|assistant|>
<think>
{round0_thinking}
</think>
‘‘‘python
{round0_operation}
<|user|>
**Environment State (Omitted)**
Previous Action Result: {round0_operation_feedback}
<|assistant|>
<think>
{round1_thinking}
</think>
‘‘‘python
{round1_operation}
<|user|>
**Environment State (Omitted)**
Previous Action Result: {round1_operation_feedback}
<|assistant|>
<think>
{round2_thinking}
</think>
‘‘‘python
{round2_operation}
<|user|>
{screenshot_for_multimodal}
* Apps: {all_apps}
* Current App: {cur_window_id}
* A11y Tree: {a11y_tree_for_text}
* App Info: {app_info}
[…]
Appendix D Training Details
During the behavior cloning stage, we construct approximately 8,000 tasks through manual annotation and data augmentation. We employ multiple advanced models to generate diverse samples for each task, and subsequently apply the eval function to filter out successful trajectories. This process yields a high-quality BC dataset comprising roughly 180,000 steps, which is then used for SFT training. We employ a 16-node computing cluster for fine-tuning, with a maximum learning rate set to $1\times 10^{-5}$, a sequence length of 32,768 tokens, and a global batch size of 256, over a total of three training epochs.
In the RL stage, the key training hyperparameters are summarized in Table 6. We initially train the BC policy (using the 1-epoch checkpoint for diversity) for 180 steps, after which performance improvements began to plateau. At this point, we collect rollouts during RL, perform task-level random selection, and curate approximately 130,000 additional steps of data for Entropulse training. The hyperparameters in this phase are identical to those used in the BC stage, except for a reduced learning rate of $5\times 10^{-6}$. RL training is then resumed until a total of 360 steps have been reached.
Table 6: Training configuration for RL training of AutoGLM-OS.
Category Parameter (Value)
Data
Task type Multi-turn chat
Max prompt length 63,488 tokens
Max response length 2,048 tokens
Train batch size 32
Responses per prompt 16
Concurrency 1024
Shuffle True
Seed 42
Actor
Exchange size $1\times 10^{10}$
Gradient checkpointing Enabled
Strategy FSDP
FSDP offloading Param + Optimizer
Sequence parallel size 2
Max tokens / GPU 32,768
Precision dtype bfloat16
Algorithm
Advantage estimator GRPO
Discount factor $\gamma$ 1.0
GAE parameter $\lambda$ 1.0
Use remove padding True
Use dynamic bsz True
Mini-batch size 32,768
Micro-batch size / GPU 1
Logprob micro-batch size / GPU 1
KL loss Enabled (low_var_kl), coef = 0.0003
Entropy coefficient 0.0
Clip ratio 0.2
Optimizer
Actor learning rate $1\times 10^{-6}$
LR warmup steps ratio 0.0
Warmup style constant
Gradient clipping 1.0
Save frequency 25
Rollout
Enable chunked prefill True
Max new tokens (generation) 2,048
Do sample True
Sampling temperature 0.8
Max turns 30
GPU memory utilization 0.7
Pools rollout 2
Pools other 6
Appendix E Human Annotation Protocol
Our annotation process involves ten trained annotators with master’s degrees, who are recruited and compensated in compliance with local labor laws and regulations. Annotators are provided with clear written guidelines to ensure consistency and accuracy, as outlined in our annotation protocol (see Figures 7 and 8). All tasks are designed to avoid sensitive personal data, and all annotated content is in English with no identifying information. The process includes task expansion—transforming generalized instructions into explicit, executable tasks—and strict result verification to minimize errors. Quality control measures include verification passes and clear formatting rules to improve annotation reliability. No annotator is exposed to harmful, discriminatory, or unsafe content during the process, and all work adheres to the Code of Ethics regarding fairness, privacy, and legal compliance.
Figure 7: Document for Human Annotation (1 / 2)
Figure 8: Document for Human Annotation (2 / 2)
Appendix F Future Direction